OpenStudy (kj4uts):

If a sample mean is 32, which of the following is most likely the range of possible values that best describes an estimate for population mean? A. (34, 42) B. (32, 42) C. (28, 36) D. (30, 38)

OpenStudy (anonymous):

Calculate the average for each of the answers and find which is closest. ie. A = (34+42)/2 = 38

OpenStudy (anonymous):

You can also see the the average is the "middle" value, so you can estimate which one has a middle value closest to 32

OpenStudy (kj4uts):

C. (28+36)/2=32

OpenStudy (anonymous):

Ya that sounds correct from my understanding of the problem :)